OpenStudy (anonymous):

Please help me i hate my life.... If 3xm+2ym-2yn-3xn=0 and "m" does not equal "n," then what is the value of y in terms of x?

OpenStudy (anonymous):

@satellite73

OpenStudy (vlery):

gimme a sec

OpenStudy (anonymous):

If we regroup the terms: 3x(m-n) + 2y(m-n)=0 (3x+2y)(m-n)=0 Since m is not equal to n, 3x+2y=0 and y= -3x/2

OpenStudy (anonymous):

Did that help? (:

OpenStudy (anonymous):

hey how come one of the m-n disappears